ABSTRACT. It is proved that if X and Y are linear spaces and F : X → P (Y) is a set-valued map with convex graph such that F (x) ≠ ϕ for all x ∈ X and F (x_{0}) is a singleton for some x_{0}, then F is single-valued and affine. Applications to metric projections and to adjoints of set-valued maps are given.